OpenStudy (brainzonly):

@Kevin Menemukan istilah 20 dan jumlah dari syarat-syarat pertama 14 kemajuan ini, 2, 5, 8, 11,...?

OpenStudy (brainzonly):

I think this is the first post in a foreign language LOL

OpenStudy (kevin):

lol, pls use english. I don't understand XD

OpenStudy (kevin):

can you translate it once more:D

OpenStudy (brainzonly):

Find the 20th term and the sum of the first 14 terms of this progression, 2, 5, 8, 11,.........?

OpenStudy (kevin):

Okay

OpenStudy (brainzonly):

U DONT speak Indonesian???????????// O.O

OpenStudy (brainzonly):

*brain explosion* XD

OpenStudy (brainzonly):

Given progression, 2, 5, 8, 11,......... is a arithmetic progression

OpenStudy (kevin):

U20 = a + (n - 1)b U20 =2 + (20 -1)3 U20 = 59

OpenStudy (kevin):

wait

OpenStudy (brainzonly):

........

OpenStudy (kevin):

Sn= 1/2n (2a + (n-1)b) S20 = 1/2*20 (2*2 + (20 -1)3) S20 = 610

OpenStudy (kevin):

am I right :D

OpenStudy (brainzonly):

Nope. LOL

OpenStudy (kevin):

:o

OpenStudy (brainzonly):

Here First term (a) = 2 and Common difference (d) = 5 - 2 = 3 We know, nth term can be represented as, an = a + (n - 1)d and sum of n terms = n2 n2 {2a + (n - 1)d} To find 20th term of the series => a20 = a + 19 d = 2 + 19 * 3 = 2 + 57 = 59 => 20th term = 59 Sum of first 14 terms = 142 142 {2 * 2 + (14 - 1) * 3} = 7{4 + 39} = 7 * 43 = 301 => Sum of first 14 terms = 301.

OpenStudy (kevin):

Owh yeah, I input it wrong. lol Go to another question :D